OpenStudy (anonymous):

why is the reciprocal for ax^2/y y/ax^2

OpenStudy (anonymous):

because they cancel so it equals 1 \[\frac{ 1 }{ 2 }\frac{ 2 }{ 1 }\] are reciprocals

OpenStudy (anonymous):

\[\frac{ ax ^{2}y }{ ax^2y }\]

OpenStudy (anonymous):

=1

OpenStudy (anonymous):

oh i think i get so is that true for all faction reciprocals? i mean does it always switch?

OpenStudy (anonymous):

a fractions reciprocal time the original fraction should equal 1

OpenStudy (anonymous):

this could help http://www.mathsteacher.com.au/year7/ch04_frac/06_recip/num.htm

OpenStudy (anonymous):

got now thank you and thanks for the link :)

OpenStudy (anonymous):

ur welcome